OpenStudy (anonymous):

Use De Moivre's theorem to express cos5* in terms of powers of sin* and cos*.

OpenStudy (anonymous):

i guess you could take \[(\cos(\theta)+i\sin(\theta))^5\] and get by demoivre \[\cos(5\theta)+i\sin(5\theta)\] and also whatever you get when you expand that thing then equate the real part to \(\cos(5\theta)\)

OpenStudy (anonymous):

i have less than no desire to multiply out \[(\cos(\theta)+i\sin(\theta))^5\] but i can't see any other way to use demoivre